# Mechanistic Understanding and Mitigation of Language Confusion in English-Centric Large Language Models

### v1

@@ Abstract
Language confusion—where large language models (LLMs) generate unintended languages against the user’s need—remains a critical challenge, especially for English-centric models. We present the first mechanistic interpretability (MI) study of language confusion, combining behavioral benchmarking with neuron-level analysis. Using the Language Confusion Benchmark (LCB), we show that confusion points (CPs)—specific positions where language switches occur—are central to this phenomenon. Through layer-wise analysis with TunedLens and targeted neuron attribution, we reveal that transition failures in the final layers drive confusion. We further demonstrate that editing a small set of critical neurons, identified via comparative analysis with multilingual-tuned models, substantially mitigates confusion without harming general competence or fluency. Our approach matches multilingual alignment in confusion reduction for most languages and yields cleaner, higher-quality outputs. These findings provide new insights into the internal dynamics of LLMs and highlight neuron-level interventions as a promising direction for robust, interpretable multilingual language modeling. Code and data will be released upon publication.
@@ 1 Introduction
Current Large Language Models (LLMs), such as GPT-4 (Achiam et al., 2023), PaLM 2 (Anil et al., 2023), and Llama 3 (Grattafiori et al., 2024), have demonstrated exceptional linguistic competence across a wide range of complex tasks that require abstract knowledge and reasoning (Dong et al., 2024; Wei et al., 2022). Early LLMs were predominantly trained on massive amounts of English text data, with some limited exposure to other languages, resulting in initially constrained multilingual capabilities (Touvron et al., 2023). Recent advances, such as multilingual continued pretraining and instruction tuning, have substantially extended these models’ ability to support multiple languages (Zhu et al., 2023; Shaham et al., 2024; Kew et al., 2024; Wang et al., 2025b). As a result, contemporary English-centric LLMs have become foundational tools for multilingual communication, multilingual content generation, and cross-lingual applications (Bang et al., 2023; Ahuja et al., 2023; Asai et al., 2024). However, despite their impressive capabilities, a persistent and underexplored limitation remains: LLMs can fail to generate text in the user’s intended language, even when explicitly instructed—a phenomenon termed language confusion (Marchisio et al., 2024). Language confusion manifests as full-response, line-level, or word-level switches into unintended languages, severely undermining user experience and model reliability, especially for non-English speakers (Figure 1a).
@@
A key observation from prior work is the identification of confusion points—specific positions in the generation process where the model abruptly switches to an unintended language. However, the model’s internal dynamics leading to these confusion points and their causal role in language confusion remain largely unexplored. This gap is particularly salient given the parallels to human bilingual code-switching, where switch points between languages are cognitively significant as extensively studied in psycholinguistics (Solorio and Liu, 2008; Bullock and Toribio, 2009). Further discussions are provided in Appendix A.
In this work, we move beyond behavioral evaluation to open the black box of LLMs, leveraging mechanistic interpretability (MI) methods (Conmy et al., 2023; Rai et al., 2024; Saphra and Wiegreffe, 2024; Sharkey et al., 2025) to investigate the internal representations and neuron-level processes underlying language confusion. We first empirically demonstrate that confusion points are critical drivers of language confusion: targeted interventions at these points can substantially reduce confusion across languages. Building on this, we employ MI tools such as TunedLens (Belrose et al., 2023) to trace the evolution of language representations through the model’s layers, revealing that confusion typically arises from transition failures in the final layers, where latent conceptual representations are mapped to surface forms in the target language (Figure 1b). To further elucidate the mechanism, we conduct a neuron-level analysis, identifying specific neurons in the last layers whose activity is predictive of successful or failed language transitions at confusion points. Inspired by recent advances in neuron attribution and editing, we show that targeted manipulation of only 100 neurons can mitigate language confusion, offering a novel, model-internal approach to improving multilingual reliability. Our findings provide the first mechanistic account of language confusion in LLMs, bridging the gap between behavioral benchmarks and internal model dynamics. By highlighting the central role of confusion points and their neural substrates, we lay the groundwork for more robust, interpretable, and cognitively informed multilingual language models.
@@
Native Human-generated es, fr, ja, ko {CJK}UTF8mj 콘크리트는 뭘로 만든거야?
@@ 2 Related Work
Mechanistic interpretability (MI) seeks to reverse-engineer neural networks by decomposing their computations into human-understandable components (Stolfo et al., 2023; Wang et al., 2024; Men et al., 2024). A central technique in MI is the projection of intermediate representations into the vocabulary space, as implemented by tools such as LogitLens (Nostalgebraist, 2020) and TunedLens (Belrose et al., 2023), which enable researchers to track how information and predictions evolve across layers (Dar et al., 2023; Pal et al., 2023). In addition to layer-wise analysis, recent work has focused on identifying, attributing, and intervening on important neurons—those whose activations are strongly correlated with specific linguistic functions or behaviors (Bau et al., 2020; Geva et al., 2022; Yu and Ananiadou, 2024b). Methods for neuron selection and editing, as well as circuit-level analysis (Elhage et al., 2021; Wang et al., 2023), have proven effective for uncovering the internal structure underlying phenomena such as factual recall (Meng et al., 2022; Geva et al., 2023), reasoning processing (Yu and Ananiadou, 2024a), and now, as in our work, language confusion. By leveraging these MI techniques, we aim to provide a granular, causal understanding of how and why language confusion arises in multilingual LLMs, and to identify actionable intervention points for mitigation.
@@
Recent research has begun to probe the internal representations of English-centric and multilingual LLMs to understand how they process and transfer information across languages (He et al., 2024; Zhao et al., 2024)Wendler et al. (2024) show that models like Llama2 often rely on English as an internal pivot language and can disentangle language and conceptual representations in controlled tasks. Fierro et al. (2025) examine how mechanisms identified in monolingual contexts generalize to multilingual settings. Wang et al. (2025a) investigate the internal causes of crosslingual factual inconsistencies, revealing how MLMs transition from language-independent to language-specific processing. However, prior work has not systematically connected these internal mechanisms to language generation errors such as language confusion.
@@ 3.1 Recap of Language Confusion Benchmark
We conducted preliminary benchmarking experiments on LCB with three instruction-tuned LLMs: Llama3-8B (English-centric, no multilingual instruction tuning), Llama3-8B-multilingual (multilingual instruction-tuned) (Devine, 2024), and Llama3.1-8B ( multilingual-optimized). As shown in Figure 1c, Llama3-8B exhibits substantial language confusion, with frequent line-level switches to unintended languages (mostly English). In contrast, both Llama3-8B-multilingual and Llama3.1-8B achieve near-perfect LPR and line-level accuracy, demonstrating the effectiveness of multilingual instruction tuning and targeted optimization for multilingual dialogue.
Given these findings, our work centers on understanding and mitigating the language confusion observed in English-centric Llama3-8B. By leveraging mechanistic interpretability methods, we aim to uncover the internal causes of confusion and develop interventions that can bring its performance closer to that of explicitly multilingual-tuned models. In the following subsection, we delve deeper into the significance of confusion points as critical junctures in the generation process.
@@ 3.2 Significance of Confusion Points
A confusion point (CP) is the position in a model’s output where the first token of an unintended language abruptly appears, marking the onset of language confusion (Marchisio et al., 2024). This concept is inspired by psycho- and neurolinguistic research on code-switching, where the precise location of a language switch—known as a switch point—is central to understanding bilingual language production and processing (Blanco-Elorrieta and Pylkkänen, 2017; Suurmeijer et al., 2020). To empirically assess the role of CPs in LLM language confusion, we conduct a replacement experiment on Llama3-8B. For each instance of language confusion, we identify the CP using the fastText language detector. We then replace the token at the CP with the corresponding token generated by Llama3-8B-multilingual, which achieves near-perfect language accuracy, under the same prompt. This approach is motivated by the psycholinguistic observation that, in human code-switching, the choice at the switch point strongly influences the subsequent language trajectory (Moreno et al., 2002; Lai and O’Brien, 2020).
@@
Our results, summarized in Table 2, show a substantial reduction in language confusion after CP replacement, even though our method does not represent an oracle upper bound. These findings highlight the centrality of confusion points in the emergence of language confusion and motivate our subsequent mechanistic analysis and targeted interventions.
@@ 4.2 Localizing Critical Neurons at Confusion Points
Figure 3: Distribution of Important Neurons Associated with Confusion Points in Llama3-8B. (a) Distribution of the top 300 most important FFN neurons across layers for an individual Chinese prompt “{CJK*}UTF8gbsn请解释拆东墙补西墙的意思。(Please explain ‘{CJK*}UTF8gbsn拆东墙补西墙.’)” from Aya. (b) Aggregated distribution of important neuron scores across all Chinese test samples in Aya.
@@ 5.1 Neuron Selection and Intervention
For each strategy, we select the top 100 neurons and intervene by setting their activations to zero during generation. We evaluate the impact of each method on the LCB benchmark. Our results (Table 4) demonstrate that Comparative Importance Selection achieves the most effective reduction in language confusion, substantially outperforming both frequency-based and aggregate importance methods. Frequency-based selection yields minimal benefit, while aggregate importance provides moderate improvement but still lags behind our proposed approach. Notably, the comparative strategy selectively targets neurons implicated in confusion, minimizing collateral impact on general model competence.
@@ 5.2 Generalization and Robustness of Neuron Editing
To further validate the effectiveness and safety of our Comparative Importance Selection strategy, we conduct a comprehensive evaluation across multiple metrics and experimental setups. Our goal is to ensure that neuron editing not only mitigates language confusion but also preserves the model’s general competence, fluency, and robustness across domains (Table 4).
@@
To evaluate whether neuron editing affects the model’s general capabilities, we test the edited model on widely used multilingual benchmarks, including XNLI and multilingual sentiment analysis. We also assess output fluency by measuring the perplexity of generated responses using the multilingual model facebook/xglm-564M (Lin et al., 2022). Across all these metrics, the edited model maintains performance comparable to the original Llama3-8B, indicating that our intervention does not degrade general language understanding or generation quality.
@@ 5.3 Comparison with Multilingual Alignment
To contextualize the effectiveness of neuron editing, we compare the performance of the edited Llama3-8B model with that of the multilingual-tuned Llama3-8B. Quantitative results show that neuron editing achieves language confusion mitigation on par with the multilingual-aligned model for most languages. Both approaches yield high line-level pass rates and language accuracy, demonstrating that targeted neuron intervention can match the benefits of extensive multilingual instruction tuning for confusion reduction.
@@ 6 Conclusions
This work provides the first mechanistic interpretability account of language confusion in English-centric LLMs. By tracing confusion points to failures in late-layer transitions and localizing the critical neurons responsible, we demonstrate that targeted neuron editing can robustly mitigate language confusion without sacrificing general competence or fluency. Our approach achieves results on par with multilingual-tuned models for most languages, while preserving cleaner output quality. These findings highlight the promise of neuron-level interventions for more reliable and interpretable multilingual language modeling.
@@ C.2 Datasets and Tasks
  • Multilingual Sentiment Analysis: Standard multilingual sentiment datasets (including German, Spanish, French, Japanese, and Chinese). It is a binary classification task derived from the multilingual Amazon review dataset.
@@ C.4 Implementation Details
All code, evaluation scripts, and neuron selection details will be released upon publication to facilitate reproducibility.

### v2

@@ Abstract
Language confusion—where large language models (LLMs) generate unintended languages against the user’s need—remains a critical challenge, especially for English-centric models. We present the first mechanistic interpretability (MI) study of language confusion, combining behavioral benchmarking with neuron-level analysis. Using the Language Confusion Benchmark (LCB), we show that confusion points (CPs)—specific positions where language switches occur—are central to this phenomenon. Through layer-wise analysis with TunedLens and targeted neuron attribution, we reveal that transition failures in the final layers drive confusion. We further demonstrate that editing a small set of critical neurons, identified via comparative analysis with a multilingual-tuned counterpart, substantially mitigates confusion while largely preserving general competence and fluency. Our approach matches multilingual alignment in confusion reduction for many languages and yields cleaner, higher-quality outputs. These findings provide new insights into the internal dynamics of LLMs and highlight neuron-level interventions as a promising direction for robust, interpretable multilingual language modeling. Code and data are available at: https://github.com/ercong21/lang_confusion.
@@ 1 Introduction
Current Large Language Models (LLMs), such as GPT-4 (Achiam et al., 2023), PaLM 2 (Anil et al., 2023), and Llama 3 (Grattafiori et al., 2024), have demonstrated exceptional linguistic competence across a wide range of complex tasks that require abstract knowledge and reasoning (Dong et al., 2024; Wei et al., 2022; Nie et al., 2025; Wang et al., 2025b). Early LLMs were predominantly trained on massive amounts of English text data, with some limited exposure to other languages, resulting in initially constrained multilingual capabilities (Touvron et al., 2023). Recent advances, such as multilingual continued pretraining and instruction tuning, have substantially extended these models’ ability to support multiple languages (Zhu et al., 2023; Tan and Monz, 2023; Shaham et al., 2024; Kew et al., 2024; Wang et al., 2025c). As a result, contemporary English-centric LLMs have become foundational tools for multilingual communication, multilingual content generation, and cross-lingual applications (Nie et al., 2023; Bang et al., 2023; Ahuja et al., 2023; Li et al., 2023; Asai et al., 2024). However, despite their impressive capabilities, a persistent and underexplored limitation remains: LLMs can fail to generate text in the user’s intended language, even when explicitly instructed—a phenomenon termed language confusion (Marchisio et al., 2024). Language confusion manifests as full-response, line-level, or word-level switches into unintended languages, severely undermining user experience and model reliability, especially for non-English speakers (Figure 1a).
@@
A key observation from prior work is the identification of confusion points—specific positions in the generation process where the model abruptly switches to an unintended language. However, the model’s internal dynamics leading to these confusion points and their causal role in language confusion remain largely unexplored. This gap is particularly salient given the parallels to human bilingual code-switching, where switch points between languages are cognitively significant as extensively studied in psycholinguistics (Solorio and Liu, 2008; Bullock and Toribio, 2009). Further discussions on the relationship between code-switching in human communication and language confusion in LLM text generation are provided in Appendix A.
In this work, we move beyond behavioral evaluation to open the black box of LLMs, leveraging mechanistic interpretability (MI) methods (Conmy et al., 2023; Rai et al., 2024; Saphra and Wiegreffe, 2024; Sharkey et al., 2025) to investigate the internal representations and neuron-level processes underlying language confusion. We first empirically demonstrate that confusion points are critical drivers of language confusion: targeted replacement at these points can substantially reduce confusion across languages. Building on the CP token, we employ MI tools such as TunedLens (Belrose et al., 2023) to trace the evolution of language representations through the model’s layers, revealing that confusion typically arises from transition failures in the final layers, where latent conceptual representations are mapped to surface forms in the target language (Figure 1b). To further elucidate the mechanism, we conduct a neuron-level analysis, identifying specific neurons in the last layers whose activity is predictive of successful or failed language transitions at confusion points. Inspired by recent advances in neuron attribution and editing, we show that targeted manipulation of only 100 neurons can effectively mitigate language confusion, offering a novel, model-internal approach to improving multilingual reliability. Our findings provide the first mechanistic account of language confusion in LLMs, bridging the gap between behavioral benchmarks and internal model dynamics. By highlighting the central role of confusion points and their neural mechanisms, we lay the groundwork for more robust, interpretable, and cognitively informed multilingual language models.
@@
Native Human-generated es, fr, ja, ko 콘크리트는 뭘로 만든거야?
@@ 2 Related Work
Mechanistic interpretability (MI) seeks to reverse-engineer neural networks by decomposing their computations into human-understandable components (Stolfo et al., 2023; Wang et al., 2024; Men et al., 2024). A central technique in MI is the projection of intermediate representations into the vocabulary space, as implemented by tools such as LogitLens (Nostalgebraist, 2020) and TunedLens (Belrose et al., 2023), which enable researchers to track how information and predictions evolve across layers (Dar et al., 2023; Pal et al., 2023). In addition to layer-wise analysis, recent work has focused on identifying, attributing, and intervening on important neurons—those whose activations are strongly correlated with specific linguistic functions or behaviors (Bau et al., 2020; Geva et al., 2022; Yu and Ananiadou, 2024b; Tan et al., 2024). Methods for neuron selection and editing, as well as circuit-level analysis (Elhage et al., 2021; Wang et al., 2023), have proven effective for uncovering the internal structure underlying phenomena such as factual recall (Meng et al., 2022; Geva et al., 2023), reasoning processing (Yu and Ananiadou, 2024a), and now, as in our work, language confusion. By leveraging these MI techniques, we aim to provide a granular, causal understanding of how and why language confusion arises in multilingual LLMs, and to identify actionable intervention points for mitigation.
@@
Recent research has begun to probe the internal representations of English-centric and multilingual LLMs to understand how they process and transfer information across languages (He et al., 2024; Zhao et al., 2024; Kojima et al., 2024; Zhang et al., 2024)Wendler et al. (2024) show that models like Llama2 often rely on English as an internal pivot language and can disentangle language and conceptual representations in controlled tasks. Fierro et al. (2025) examine how mechanisms identified in monolingual contexts generalize to multilingual settings. Wang et al. (2025a) investigate the internal causes of crosslingual factual inconsistencies, revealing how multilingual models transition from language-independent to language-specific processing. However, prior work has not systematically connected these internal mechanisms to language generation errors such as language confusion.
@@ 3.1 Recap of Language Confusion Benchmark
We conducted preliminary benchmarking experiments on LCB with three instruction-tuned LLMs: Llama3-8B (English-centric, no multilingual instruction tuning), Llama3-8B-multilingual (multilingually instruction-tuned) (Devine, 2024), and Llama3.1-8B (multilingually optimized). As shown in Figure 1c, Llama3-8B exhibits substantial language confusion, with frequent line-level switches to unintended languages (mostly English). In contrast, both Llama3-8B-multilingual and Llama3.1-8B achieve near-perfect line-level accuracy, demonstrating the effectiveness of multilingual instruction tuning and targeted optimization for multilingual dialogue.
Given these findings, our work centers on understanding and mitigating the language confusion observed in English-centric Llama3-8B. By leveraging mechanistic interpretability methods, we aim to uncover the internal causes of confusion and develop interventions that can bring its performance closer to that of explicitly multilingually-tuned models. In the following subsection, we delve deeper into the significance of confusion points as critical junctures in the generation process.
@@ 3.2 Significance of Confusion Points
A confusion point (CP) is the position in a model’s output where the first token of an unintended language abruptly appears, marking the onset of language confusion (Marchisio et al., 2024). This concept is inspired by psycho- and neurolinguistic research on code-switching, where the precise location of a language switch—known as a switch point—is central to understanding bilingual language production and processing (Blanco-Elorrieta and Pylkkänen, 2017; Suurmeijer et al., 2020). To empirically assess the role of CPs in LLM language confusion, we conduct a replacement experiment on Llama3-8B. For each instance of language confusion, we identify the CP using the fastText language detector. We then replace the token at the subword-level CP with the corresponding token generated by Llama3-8B-multilingual, which achieves near-perfect language accuracy, under the same prompt. This approach is motivated by the psycholinguistic observation that, in human code-switching, the choice at the switch point strongly influences the subsequent language trajectory (Moreno et al., 2002; Lai and O’Brien, 2020).
@@
Our results, summarized in Table 2, show a substantial reduction in language confusion after CP replacement, even though our method does not represent an oracle upper bound. These findings highlight the centrality of confusion points in the emergence of language confusion and motivate us to base the subsequent mechanistic analysis and targeted interventions on the generation of the token at the CP position.
@@ 4.2 Localizing Critical Neurons at Confusion Points
Figure 3: Distribution of Important Neurons Associated with Confusion Points in Llama3-8B. (a) Distribution of the top 300 most important FFN neurons across layers for an individual Chinese prompt “请解释拆东墙补西墙的意思。(Please explain ‘拆东墙补西墙.’)” from Aya. (b) Aggregated distribution of important neuron scores across all Chinese test samples in Aya.
@@ 5.1 Neuron Selection and Intervention
For each strategy, we select the top 100 neurons and intervene by setting their activations to zero during generation. We evaluate the impact of each method on the LCB benchmark. Our results (Table 3) demonstrate that Comparative Importance Selection achieves the most effective reduction in language confusion, substantially outperforming both frequency-based and aggregate importance methods. Frequency-based selection yields minimal benefit, while aggregate importance provides moderate improvement but still lags behind our proposed comparative approach. Notably, the comparative strategy selectively targets neurons implicated in confusion, minimizing collateral impact on general model competence.
@@ 5.2 Generalization and Robustness of Neuron Editing
To further validate the effectiveness and safety of our Comparative Importance Selection strategy, we conduct a comprehensive evaluation across multiple metrics and experimental setups. Our goal is to ensure that neuron editing not only mitigates language confusion but also preserves the model’s general competence, fluency, and robustness across domains (Table 4). Details of the experimental setup for the generalization and robustness experiments are provided in Appendix C.
@@
To evaluate whether neuron editing affects the model’s general capabilities, we test the edited model on widely used multilingual benchmarks, including XNLI (Conneau et al., 2018) and multilingual sentiment analysis (Keung et al., 2020). We also assess output fluency by measuring the perplexity of generated responses using the multilingual model facebook/xglm-564M (Lin et al., 2022). Across all these metrics, the edited model maintains performance comparable to the original Llama3-8B, indicating that our neuron intervention method does not degrade general language understanding or generation quality.
@@ 5.3 Comparison with Multilingual Alignment
To contextualize the effectiveness of neuron editing, we compare the performance of the edited Llama3-8B model with that of the multilingual-tuned Llama3-8B. Quantitative results show that neuron editing yields confusion mitigation comparable to multilingual instruction tuning for many languages, despite requiring no additional data or retraining. This demonstrates that targeted neuron intervention can match the benefits of extensive multilingual instruction tuning for confusion reduction. However, for languages with limited pretraining exposure, multilingual supervised fine-tuning can achieve stronger performance.
@@ 6 Conclusions
This work provides the first mechanistic interpretability account of language confusion in English-centric LLMs. By tracing confusion points to failures in late-layer transitions and localizing the critical neurons responsible, we demonstrate that targeted neuron editing can robustly mitigate language confusion without sacrificing general competence or fluency. Our approach achieves results comparable with multilingual-tuned models for many languages, while preserving cleaner output quality, though multilingual SFT remains stronger for some low-resource cases. These findings highlight the promise of neuron-level interventions for more reliable and interpretable multilingual language modeling.
@@
Acknowledgments
We want to thank the anonymous reviewers for their valuable feedback. This work was supported by the German Research Foundation (DFG) under grant SCHU 2246/14-1, Munich Center for Machine Learning (MCML), and China Scholarship Council (CSC).
@@ C.2 Datasets and Tasks
  • Multilingual Sentiment Analysis: Standard multilingual sentiment datasets (including German, Spanish, French, Japanese, and Chinese). It is a binary classification task derived from the multilingual Amazon review dataset (Keung et al., 2020).